A construction worker needs to raise some materials to the top of a 20-ft tall building by using a computer-controlled crane. If the materials are 28 feet away from the base of the building, what slope needs to be programmed into the computer to accomplish the delivery? State your answer as a fraction in lowest terms


1
Expert's answer
2021-10-14T08:03:43-0400

If the information above would be presented on a graph then,

  1. The coordinates of the top of the building would be "(x2,y2)=(0,20)"
  2. The coordinates of the point where the materials are is "(x1,y1)=(28,0)"

If a line is drawn to connect the two points then, the slope/gradient "(m)" of this line is what we find in order for it to be programmed into the computer to accomplish the delivery.

To find slope "(m)" , the formula below is used.

"m=(y2-y1)\/(x2-x1)"

Given the values of "(x1,y1)=(28,0)" and "(x2,y2)=(0,20)" then the slope"(m)" is given as,

"m=(20-0)\/(0-28)=20\/(-28)=-(20\/28)=-5\/7" .

Therefore, the slope that needs to be programmed into the computer to accomplish the delivery is "m=-5\/7"